### v2

# Challenges and Requirements for Benchmarking Time Series Foundation Models

Abstract
Time Series Foundation Models (TSFMs) represent a new paradigm for time-series forecasting, promising zero-shot predictions without the need for task-specific training or fine-tuning. However, similar to Large Language Models (LLMs), the evaluation of TSFMs is challenging: as training corpora grow increasingly large, it becomes difficult to ensure the integrity of the test sets used for benchmarking. Our investigation of existing TSFM evaluation studies identifies two kinds of information leakage: (1) train-test sample overlaps arising from the multi-purpose reuse of datasets and (2) temporal overlap of correlated train and test series. Ignoring these forms of information leakage when benchmarking TSFMs risks producing overly optimistic performance estimates that fail to generalize to real-world settings. We therefore argue for the development of novel evaluation methodologies that avoid pitfalls already observed in both LLM and classical time-series benchmarking, and we call on the research community to adopt principled approaches to safeguard the integrity of TSFM evaluation.
1 Introduction
Time Series Foundation Models (TSFMs) represent an emerging paradigm in forecasting, drawing inspiration from the architecture and training methodologies of foundation models in natural language processing (NLP). Despite being a relatively new research area, the field has evolved into a highly dynamic landscape featuring a wide variety of architectures and approaches. This rapid proliferation is exemplified by diverse models such as Chronos (Ansari et al., 2024a), Chronos-bolt (Ansari et al., 2024b), Chronos-2 (Ansari et al., 2025), Moirai/Moirai-MoE (Woo et al., 2024; Liu et al., 2024b), MOMENT (Goswami et al., 2024), Time-MoE (Shi et al., 2024), TimesFM (Das et al., 2024), TimesFM-2.0 (Das et al., 2024), TimesFM-2.5 (Das et al., 2024), or TiRex (Auer et al., 2025). In contrast to traditional time series models that are trained from scratch for specific tasks and time series, TSFMs leverage transfer learning: they are pre-trained on large collections of generic time series and can perform zero-shot forecasting without domain adaptation or fine-tuning (Liang et al., 2024).
The reliance on massive amounts of time series training data, typically scraped from public repositories, raises the question of whether the proper evaluation of TSFMs faces similar challenges to those observed in the evaluation of Large Language Models (LLMs). In the NLP domain, training on vast portions of the internet has given rise to an “evaluation crisis" (Liao and Xiao, 2023), in which test set contamination (Mirzadeh et al., 2024; Ravaut et al., 2024; Li et al., 2024a) and memorization effects (Chang et al., 2024) have led to overly optimistic performance estimates. Whether current TSFM evaluations are vulnerable to analogous issues remains an open question, with important implications for fair and reliable model comparison.
While benchmark strategies such as benchmarking on held-out data (Aksu et al., 2024) and clean train-test splits (Qiu et al., 2024) have been proposed recently, we doubt that these approaches effectively prevent information leakage in the context of TSFMs. Beyond the risk of direct information leakage, the increased memorization capacity of scaled foundation models poses additional challenges (see Section 3).
This article examines the validity and reliability of current TSFM evaluation practices and discusses the risk of an evaluation crisis in the TSFM domain. We address this issue through two contributions: first, we conduct a comprehensive train-test set lineage analysis of 22 leading TSFMs and review all major evaluation benchmarks available as of early 2026. Second, we empirically demonstrate the phenomenon of temporal overlaps of correlated train and test series, a challenge specific to TSFMs, through a dedicated experiment.
Our investigation yields two primary findings that suggest that TSFM evaluation is indeed compromised:
  • First, while we find no evidence of unintended direct information leakage within the training and evaluation of individual TSFMs, our analysis reveals substantial ambiguity in data separation between models. In several cases, one model’s training set serves as another’s test set. This issue is further exacerbated by dataset remixing, renaming, and inadequate documentation, which already pose significant evaluation challenges for newer models (see Section 3.1).
  • Second, we identify the temporal overlap of correlated train and test series as an indirect form of information leakage specific to TSFMs. External shocks, such as the COVID-19 pandemic to stock prices, as well as exogenous drivers like local weather conditions to local solar production, act as factors that directly shape or distort the causal dynamics between multiple time series (see Section 3.2). As a result, a TSFM trained and evaluated on disjoint time series covering the same time period can effectively memorize patterns induced by such causal relations (see Figure 3).
Based on these findings, we argue that the rise of TSFMs necessitates new benchmarking methodologies to prevent an evaluation crisis akin to that observed in NLP.
2 Background
2.1 Time Series Foundation Models
A foundation model is defined as “any model that is trained on broad data (generally using self-supervision at scale) that can be adapted or fine-tuned to a wide range of downstream tasks” (Bommasani et al., 2021). The term is primarily associated with Transformer-based Large Language Models (LLMs) like the ChatGPT or Llama families (OpenAI, 2023; Grattafiori et al., 2024). Their success is largely attributed to scaling laws (Kaplan et al., 2020), describing the practice of increasing training data, parameters, and compute to improve a model’s predictive performance. Although LLMs demonstrate impressive zero-shot capabilities across benchmarks (Liang et al., 2022), concerns persist that they partly rely on memorizing benchmark data (Chang et al., 2024; Carlini et al., 2022). Such test set contamination, where test samples appear in pre-training data like Common Crawl [^1] https://commoncrawl.org/ or The Pile [^2] https://pile.eleuther.ai/, remains common despite improved dataset curation, leading to inflated performance estimates (Li et al., 2024a; Singh et al., 2024; Grattafiori et al., 2024; Mirzadeh et al., 2024; Ravaut et al., 2024).
Recognizing the structural similarity between language modeling and time series analysis, recent research has adapted the Transformer architecture to forecast future values of time series, ranging from the classical encoder-decoder Transformer (Ansari et al., 2024a) over Decoder-only models (Das et al., 2024) to Encoder-only models often applied in multivariate time series forecasting (Goswami et al., 2024). Beyond this, TSFMs are based on a variety of other architectural backbones such as LLMs Gruver et al. (2024); Xue and Salim (2024), Vision-based Autoencoders Chen et al. (2024), and mixed architectures that leverage LLM reprogramming Jin et al. (2024). Across these approaches, a key commonality lies in how data are represented: LLM-based models encode numerical values directly as language tokens like LLMTime (Gruver et al., 2024) and PromptCast (Xue and Salim, 2024), whereas some Transformer-based variants rely on specialized tokenization schemes to discretize continuous time series, as exemplified by TimeGPT-1 (Garza et al., 2024) and the T5-based Chronos (Ansari et al., 2024a). Furthermore, Vision-based approaches represent time series as gray-scaled images, while reprogramming-based approaches mix textual and continuous time series representations Jin et al. (2024). Similar to LLMs, all presented architectures are pre-trained on large collections of generic time series and appear to follow scaling laws (Edwards et al., 2024), enabling them to produce zero-shot forecasts and handle a wide range of frequencies and forecasting horizons. Only one exception—MOMENT—requires task-adaptive fine-tuning for high-quality forecasts Goswami et al. (2024). Based on an encoder-only architecture this model is capable of time series regression, classification, anomaly detection, and forecasting Goswami et al. (2024). However, due to its broad application area, the model needs task-adaptive fine-tuning of the last neural network layer that is specialized for each task. A key distinction between the foundation models in NLP and time series analysis concerns the definition of “zero-shot” capabilities. While LLMs typically adapt to entirely different tasks (Brown et al., 2020), TSFMs typically perform zero-shot forecasting by generalizing across domains, time series, input lengths, frequencies, and horizons without the need for model retraining (Ansari et al., 2024a; Das et al., 2024; Woo et al., 2024).
2.2 Benchmarking Methodologies
Traditional statistical or machine-learning-based models for time series forecasting generally require explicit training on the target time series. These local models exhibit only limited generalization capabilities; that is, they can only predict values that are identically distributed to their training data. Consequently, the dominant logic in time series benchmarking relies on time-based train/test splits (Hyndman and Athanasopoulos, 2021). This approach creates a temporal separation between training and evaluation data to prevent information leakage and is employed in major time series collections such as the Monash repository (Godahewa et al., 2021) and TFB (Qiu et al., 2024).
To rigorously assess a model’s predictive capabilities under evolving temporal dynamics, Time-Series Cross-Validation (TSCV) is widely regarded as best practice. Instead of using a single time-based split, TSCV uses the idea of rolling windows (see left in Figure 1). This evaluation strategy does not necessarily require retraining; instead, the training set can be fixed while the inference window is rolled forward, reducing evaluation costs to inference-only computation (Hewamalage et al., 2023).
Forecasting competitions, like the M-competitions or Kaggle, try to avoid the risk of test set contamination by soliciting forecasts for unpublished private test sets (Makridakis et al., 2022; Bojer and Meldgaard, 2021). These competitions are typically held at irregular intervals.
While all of the above evaluation strategies rely on historical data for testing, the recent M6 competition introduced a live evaluation using future values of financial assets on selected dates as the target (Makridakis et al., 2024).
Figure 1: Evaluation strategies along the time (left) or the domain dimension (right).
While current TSFM benchmarking studies also rely on historical evaluation data, their logic diverges from the above described time-series cross-validation strategy. Instead of using TSCV on one series as illustrated in Figure 1 (a), they typically evaluate models using many series but with only a single time-based split per series (Das et al., 2024; Gruver et al., 2024). Figure 1 (b) shows the focus of most TSFM evaluations. Here, the focus is not on temporal depth, but on generalization across heterogeneous series, especially in zero-shot scenarios where complete datasets or domains are held out of the pre-training (Shi et al., 2024; Ansari et al., 2024a). This approach is focused on the generalization capabilities of TSFMs and prevents cherry-picking of test sets. Two prominent examples of this strategy are TSFM-Bench, which comprises 21 test sets with time series from diverse domains (e.g., energy, finance), statistical characteristics (e.g., stationary, trend), and frequencies (e.g., quarter-hourly, hourly) Li et al. (2025) and GIFT-Eval, which comprises 23 curated test sets with time series of seven domains, ten frequencies, and multivariate inputs (Aksu et al., 2024).
While this series cross validation focuses on evaluating the generalization capabilities of TSFMs, it is not without drawbacks. Roque et al. (2025) showed that with just four cherry-picked test sets, out of the distribution of available test sets, 46% of models could appear as state-of-the-art purely due to selection bias. Notably, deep learning models were found to be more susceptible to this variance than classical baselines. Especially when evaluation general-purpose TSFMs, generalization is a core evaluation focus.
Figure 1 (c) illustrates the current best practice in TSFM evaluation, that is, a combined time and series cross validation (Yu et al., 2016; Salinas et al., 2020). While this approach has first adopters, notably Goktas et al. (2025b); Shchur et al. (2025), it is not yet commonly used in TSFM benchmarking.
The scaling of pre-training and evaluation datasets in TSFMs gives rise to another challenge that can also be observed in LLM evaluation. As LLMs are pre-trained on large datasets crawled from the internet like Common Crawl [^3] https://commoncrawl.org/ or The Pile [^4] https://pile.eleuther.ai/, they have already seen most benchmark tasks during pre-training. Ultimately, this led to an LLM “evaluation crisis” Liao and Xiao (2023). Likewise, with large and obscure pre-training datasets, defining test datasets that do not overlap with models’ pre-training datasets becomes increasingly difficult in TSFMs Aksu et al. (2024); Goktas et al. (2025a); Xu et al. (2025). As shown in the GIFT-Eval ablation studies, information leakage can strongly inflate evaluation performance of TSFMs Aksu et al. (2024). Thus, investigating information leakage in TSFMs is an important research direction. Addressing this, current benchmarks developed strategies to prevent information leakage. GIFT-Eval provides an explicit pretraining dataset containing approximately 230 billion data points and a dedicated test with 177 million data points Aksu et al. (2024) staying with a series cross validation (see Figure 1 (a)) but no time cross validation. Two other benchmarks TempusBench and Fidel-TS, use test datasets from unknown or restricted sources to prevent information leakage Goktas et al. (2025a); Xu et al. (2025). Particularly, TempusBench leverages a new test set with 48 uni- and multi-variate time series forecasting tasks that is currently not in any TSFM pre-training corpus Goktas et al. (2025a). While they apply a time and series cross-validation strategy, they further propose to leverage real-world live data for evaluations Goktas et al. (2025a). Fidel-TS apply a similar strategy by leveraging restricted live APIs but implement a series cross validation without time cross validation Xu et al. (2025).
Overall, all current approaches to mitigate information leakage and the possibly resulting inflation of performance rely on the development of new datasets that can currently not be in the pre-training corpus of TSFMs.
3 Challenges of Evaluating Time Series Foundation Models
Based on the need for an examination of information leakage in current TSFMs, we decided to first conduct a TSFM train-test dataset lineage analysis (Section 3.1). We then further analyze the outcomes of potential indirect overlaps in time series between pre-training and test sets of TSFMs.
3.1 Direct Information Leakage: Train-Test Sample Overlap
To examine potential test set contamination in existing TSFM evaluations, we searched for peer-reviewed and accepted papers including the terms "Time Series Foundation Model" at leading venues in the machine learning area (TMLR, ICML, NeurIPS, ICLR, WWW) and their workshops, published up to January 2026. This search resulted in 22 published TSFMs that met our inclusion criteria (Table 1). For each of these papers, we analyzed their dataset lineage, categorizing the used datasets into pre-training, train/test and zero-shot.
Table 1: Time series foundation models published in recent years
Model Name Author Year Conference/Journal
ForecastPFN Dooley et al. (2023) 2023 NeurIPS 23
GPT4TS Zhou et al. (2023) 2023 NeurIPS 23
Lag-Llama Rasul et al. (2024) 2024 NeurIPS 23 Workshop R0-FoMo
LLMTime Gruver et al. (2024) 2024 NeurIPS 23
Time-LLM Jin et al. (2024) 2024 ICLR 24
Chronos Ansari et al. (2024a) 2024 TMLR
TimesFM Das et al. (2024) 2024 ICML 24
Moirai Woo et al. (2024) 2024 ICML 24
MOMENT Goswami et al. (2024) 2024 ICML 24
Timer Liu et al. (2024c) 2024 ICML 24
UniTime Liu et al. (2024a) 2024 WWW 24
TinyTimeMixer Ekambaram et al. (2024) 2024 NeurIPS 24
Time-MoE Shi et al. (2024) 2025 ICLR 25
Moirai-MoE Liu et al. (2024b) 2024 ICML 25
VisionTS Chen et al. (2024) 2024 ICML 25
LightGTS Wang et al. (2025a) 2025 ICML 25
Sundial Liu et al. (2025) 2025 ICML 25
TimesFM-IC Faw et al. (2025) 2025 ICML 25
ROSE Wang et al. (2025b) 2025 ICML 25
FlowState Graf et al. (2025) 2025 NeurIPS 25 Workshop BERT2S
TiRex Auer et al. (2025) 2025 NeurIPS 25
Toto Cohen et al. (2025) 2025 NeurIPS 25
Pre-training datasets are used to fit a model to general time series patterns, such as, trends and seasonality. They are employed for initial training only and, hence, should be excluded from any evaluation activities. Train/test datasets, also sometime called in-domain or in-distribution datasets in the context of TSFMs, are datasets used to fine-tune a model to domain-specific patterns. They are split at a defined point in time into training and test sets and the model is fitted on the training data and evaluated on the test data (Hyndman and Athanasopoulos, 2021; Kapoor and Narayanan, 2023). Zero-shot datasets are used for evaluation of TSFMs on unseen data and should therefore be excluded from any training and fine-tuning. In a zero-shot setting, a model is given a short historical context of an unseen time series and it is assessed how well it can predict future values of this series.
Figure 2: Lineage of dataset collections (left) used for training and evaluating recent Time Series Foundation Models (right). Typically, a collection contains multiple different datasets. Lines indicate cases where at least one dataset of a collection was used for pre-training, train/test, or zero-shot evaluation. Smaller datasets (used n<10) are shown as "Other".
Our lineage analysis, the details are described in Supplementary Information Section D and the results are visualized in Figure 2, reveals that there is no consensus between model developers which publicly available datasets should be used for what purpose. In fact, each TSFM team uses a different combination of pre-training, train/test, and zero-shot datasets.
Whereas we find no indication that any individual TSFM suffers from information leakage within its own training and evaluation, the widespread practice of re-purposing datasets introduces substantial challenges for benchmarking across TSFMs. Once a dataset has been used for pre-training or fine-tuning by any model, it must be regarded as contaminated and is no longer admissible for TSFM benchmarking. Our lineage analysis shows that this is not merely a theoretical concern, but a concrete and pervasive practical issue. No dataset has been universally used across all TSFMs for train/test or zero-shot evaluation (see Supplementary Information Section D). And only 6% of the analyzed datasets have never been used for pre-training or fine-tuning, making them the only suitable candidates for genuine zero-shot model benchmarking.
In Supplementary Information Section B, we summarize literature showing that leakage between training and test sets can easily lead to over 50% better test scores. Several documented leakage cases suggest that the often unintended overlaps between train and test data are not always identified during the peer-review process Saravanan et al. (2024); Montet et al. (2025); Li et al. (2024b). This is not surprising as exactly determining on which datasets a model has been trained requires careful reading of the original paper, its appendix, and sometimes even analyzing published source code. Doing this for a dozen or more models when performing a benchmarking study requires substantial effort from the authors and reviewers as well. As more TSFMs are published, authors increasingly face the challenge of navigating complicated dataset relationships. For example, the authors of the Tirex model discovered that a dataset called "Solar" appears in parts in the Gift-Eval test data Aksu et al. (2024) as well as in the Chronos pre-training data, which is used by several models including Tirex, Toto, and Sundial Ansari et al. (2024a); Auer et al. (2025); Liu et al. (2025); Cohen et al. (2025). As a precaution, Tirex and Toto excluded the dataset, though Tirex explicitly notes the difficulty of verifying whether the data is actually identical. Through our lineage analysis, we found that in the Chronos pre-training data, the "Solar" dataset includes solar power data from all US states at two frequencies: 5T (5 minutes) and 1H (hourly), where the 5T data was sourced directly from the original data provider [^5] https://www.nlr.gov/grid/solar-power-data. Gift-Eval, however, sourced only the Alabama subset as a test set, aggregated to 10T, 1H, 1D (daily), and 1W (weekly), through the LSTNet dataset repository [^6] https://github.com/laiguokun/multivariate-time-series-data/tree/master/solar-energy. LSTNet also sourced the data from the original provider but, as noted, included only Alabama rather than all states. In addition, the Alabama solar subset appears in the Monash Repository [^7] https://forecastingdata.org/ at frequencies 10T and 1W. This modified version of the source data was used as pretraining data by the models TinyTimeMixers and Lag-Llama Ekambaram et al. (2024); Rasul et al. (2024). While TinyTimeMixers and Lag-Llama were published before Gift-Eval, the authors of Sundial validate against Gift-Eval and state they excluded any overlaps without explicitly listing which datasets they excluded. This example illustrates how dataset provenance can become convoluted through multiple transformations and repositories, making it difficult for researchers to identify overlaps.
While in the previous example the dataset name was constantly "Solar", relying on dataset names to avoid train-test overlaps is not sufficient. The identification of suitable datasets is complicated by dataset remixing and renaming. For example, the dataset "Elecdemand" from the Monash Repository is a scaled subset ($1/1000$) of the “Australian Electricity Demand” dataset. Similarly, the dataset "ElectricityLoadDiagrams20112014" appears under various names, such as "Electricity" in the Autoformer and Monash collections (Godahewa et al., 2021; Wu et al., 2022) and "ECL" in the Informer (Zhou et al., 2021) collection. Dataset collections sometimes also contain different observations of the the same variables. For instance, the car ride shares dataset is included with same frequency and from the same year 2015 in both the Monash and GluonTS collections. Although both repositories cite the same GitHub source [^8] https://github.com/fivethirtyeight/uber-tlc-foil-response, a close examination reveals that they contain different time series of distinct For-Hire Vehicles (Uber in GluonTS vs Lyft in Monash). Consequently it is necessary to always compare the underlying data itself, keeping in mind it could have been transformed in various ways (scaling, handling missing values, renaming, resampling, etc.).
3.2 Indirect Information Leakage: Temporal Overlap of Correlated Train and Test Series
Direct overlap in data points between training and test sets is not the only cause for information leakage. In the context of time series data, a more subtle case of information leakage can happen due to indirect dependencies between data points in the training and test sets (Kapoor and Narayanan, 2023). Often, these indirect dependencies stem from confounding bias, that is, an exogenous factor influencing observations from both datasets. For example, global events or crises like COVID-19 or the European energy crisis, caused by Russia’s invasion of Ukraine, can influence a large number of time series from different domains. However, not only global phenomena but also local factors can introduce such dependencies. Weather conditions at a specific location, for instance, directly influence solar energy production data, creating a causal relationship between meteorological measurements and power generation time series. Similarly, local traffic patterns may affect both air quality measurements and delivery time predictions within the same geographic area.
A key strength of TSFMs lies in their ability to learn subtle patterns in a flexible, data-driven manner and to apply these patterns to forecast future values of previously unseen time series. At large scale, however, TSFMs may inadvertently memorize patterns observed in one or more training time series and transfer them to series in the test set from the same time period. As long as training and test series are correctly ordered in time and do not overlap, this behavior does not constitute a serious issue. However, if training and test series overlap temporally and have a casual relationship (confounding or direct causation), such pattern transfer can result in information leakage. In the example illustrated in Figure 3, a highly flexible TSFM may learn the characteristic shape of the stock market decline caused by the global COVID-19 pandemic from the time series of the DAX index and reproduce it when tasked with forecasting the S&P 500 index over the same period.
Figure 3: Indirect Information Leakage - Temporal Overlap of Correlated Train Series A and Test Series B.
This temporal overlap of correlated train-test samples introduces a problematic dependency between the training and test data, violating the assumption of independence that is crucial for robust evaluation. As a result, the model gains access to information during training that would be inaccessible in a real-world scenario at the time of prediction, leading to inflated performance metrics. Therefore the explicit time period of a pattern and its actually encoded information gain importance for globally trained models such as TSFMs.
This issue is not merely theoretical, but has been demonstrated empirically. Rodrigo and Ortiz (2024) conducted an experiment isolating the influence of temporal overlap information leakage. They build a setup with highly correlated time series, such as metro, bus, road, and train usage in Madrid for the year 2024. They trained two models in a global manner with all series. One model was trained with a common train-test split date for all series. For the other model only the metro time-series was cutted at the train-test split date, keeping the other series fully in the train set. After training, they only predicted the test period of the metro series, which was never part of the training dataset. The second model with full access to the other correlated series during training achieved a Mean Absolute Error (MAE) of 248,24 compared to 439,14 to the first model with a common train-test split date across all series. This corresponds to a performance improvement of approximately 43.47% attributable solely to this indirect form of leakage, confirming that access to correlated and temporally overlapping external information significantly distorts evaluation results even without direct access to the target series.
Figure 4: Zero-shot forecasting for S&P 500 stock index during the Covid shock in the year 2020. In blue is the ground truth, in green a pretrained model with data up to December 2019, in red a model pretrained until December 2020. In both cases S&P 500 is not part of the training data.
While the study by Rodrigo and Ortiz (2024) used time series where the complete training data is highly correlated and temporally overlapping with the complete test data, we replicate an TSFM training process in a more realistic scenario. We incorporate diverse data where only a small percentage of observations is affected by indirect leakage. To demonstrate the impact of temporal train-test leakage on model performance under these conditions, we conducted an experiment using a trained-from-scratch Time-Series Transformer (TST) (Zerveas et al., 2020). As we have seen in Section 3.1, any other pre-trained TSFM could be already affected by this form of information leakage. The encoder-only model shares multi-variate characteristics with recent models such as MOMENT or Chronos-2 (Goswami et al., 2024; Ansari et al., 2025) but is significantly smaller, enabling faster training saturation. We trained the model using 10 different random seeds for two different training sets [^9] Details on the general setup, used datasets, scaling, hyperparameter-tuning and results can be found in Supplementary Information Section C: one with a temporal overlap to the test set during the COVID-19 pandemic and one without such overlap. The first training set comprised the NN5 Daily and Tourism Monthly (Monash collection), and the American (Dow Jones), Brazilian (Bovespa), British (FTSE 100), Canadian (TSX), French (CAC 40), German (DAX) and Japanese (Nikkei 225) stock indices from 2015 to the end of 2019. The only difference in the second training set was that the stock index data spanned from 2016 to the end of 2020, comprising observations affected by the COVID-19 outbreak. No data points were shared directly between the training and test sets.
Table 2: Per-seed S&P 500 MAE breakdown for indirect leak from other countries stock indices vs no-leak experiments. The percentage difference shows how much better the model with leakage data performs than the non-leakage model. For each seed, the better MAE in displayed in bold.
Seed MAE Before Split MAE After Split
No Leak Indirect Leak % Diff No Leak Indirect Leak % Diff
410 230.53 227.72 1.22% 157.85 199.33 -26.28%
1425 219.68 178.09 18.93% 120.77 82.53 31.67%
1680 193.52 187.19 3.27% 80.74 81.22 -0.60%
1825 247.76 156.12 36.99% 110.47 70.36 36.31%
2287 231.19 206.28 10.78% 102.42 100.54 1.84%
3658 219.23 156.91 28.43% 117.25 84.62 27.84%
4013 224.75 193.16 14.06% 114.19 97.98 14.20%
4507 206.45 213.70 -3.51% 93.10 84.01 9.77%
8936 189.40 199.38 -5.27% 103.59 153.67 -48.35%
9675 206.30 182.99 11.30% 82.35 132.44 -60.82%
Mean 216.88 190.15 11.62% 108.27 108.67 -1.44%
Std Dev 17.15 21.88 21.10 38.83
We selected stock prices as our target variable as they are specifically hard to model appropriately in a long-term valid manner Goyal et al. (2024), suggesting that any performance gaps must come from external leaked information. For both models, we used the S&P 500 index as a zero-shot test set, which was not part of both datasets. Additionally, we chose a time-frame where a common cause had influence on many different time series and chose therefore the COVID-19 outbreak. In Figure 3, time series 1 shows the DAX index and time series 2 represents the S&P 500 index. Note how the COVID-19 outbreak is visible in both series.
Figure 4 displays the mean and standard deviation of the predictions for both models across the ten runs, and Table 2 shows the specific values for each seed. For almost all seeds, the model trained on data with temporal overlap to the test set predicted the initial market crash following the outbreak significantly better than the non-leakage model. In total, the temporal overlap model performed better eight out of ten times. Up to the first of May, the leakage model achieved an average MAE of about 190 compared to about 217 for the non-leakage model. For the remaining part, both models had a very similar average MAE, with the non-leakage model being slightly better (108.3 compared to 108.7 for the non-leakage model. When looking at the specific runs, shown in Table 2, the performance difference between the two models is up to about 37% in favor of the leakage model during the shock. Despite the temporal overlap representing only a small fraction of the overall training data, specifically seven time series out of 484 (or 1,600 out of 205,859 observations in total) (Table 6), this performance gap underscores the significant impact the small temporal overlap of correlated series can have on the predictive performance of a model.
4 TSFM Benchmark Requirements
The essence of time series forecasting is to predict the future. Paradoxically, using data from the past to evaluate their ability to do so is standard practice (Section 2.2). Before the advent of TSFMs, this paradox was not a serious issue, as traditional time series models have to be trained on past values of the same time series they are expected to continue. The only pitfall to avoid is adhering to the correct ordering of training and test observations.
However, the paradigm shift behind TSFMs introduces new possibilities for information leakage. Unlike local models, TSFMs learn patterns globally. Consequently, strict temporal ordering within a single time series is no longer sufficient to guarantee a fair evaluation. To rigorously assess these models, we must address leakage across two distinct dimensions: the sample dimension and the temporal dimension (see Figure 5).
Figure 5: Information leakage risks and challenges arising from the advent of TSFMs, and the requirements needed to address them.
Flexibility in Pre-Training Data (Training Dimension) The current lack of transparency surrounding TSFM pre-training corpora creates a substantial risk of information leakage as shown in Section 3.1. This growing confusion is evident in newer benchmarks and models that report results contaminated by data leakage—whether unintentional or intentional Aksu et al. (2024); Auer et al. (2025); Li et al. (2024b). Attempting to mitigate this risk by restricting pre-training data is both counterproductive and ineffective: such restrictions artificially limit the potential of foundation models and, given the opacity of large-scale pre-training corpora, our lineage analysis shows that it is almost impossible to rule out train-test sample overlaps. We therefore argue for full flexibility in pre-training data. A robust evaluation framework should be agnostic to pre-training data choices, allowing the use of any existing data to fully exploit the predictive capabilities of TSFMs. This perspective shifts the responsibility for preventing leakage from the TSFM developer to the benchmark designer, thereby necessitating evaluation datasets that are genuinely novel and capable of assessing generalization rather than memorization.
Strict Global Post-Training Test Periods (Inference Dimension) Indirect temporal dependencies between time series, originating from direct or indirect causal relationships, create another more subtle risk for information leakage. Preventing this risk by enforcing global cut-off dates for pre-training data to separate training from test data is not a viable solution to this challenge. First, it would require all researchers to tailor their training protocol to exactly this date and second, it prevents new time series patterns of current time periods after the cut-off from being trained into the models. We therefore argue for strict global post-training test periods. These test periods need to be flexibly determined based on the global last pre-training time point of all evaluated models, instead of forcing a pre-training cut-off. For a robust evaluation framework a strict temporal barrier must be enforced across the entire universe of data: if the last pre-training timestamp across all evaluated models is $t$, only observations from $t+1$ onwards can be present in the test set of a benchmark, ensuring that TSFMs rely solely on past dynamics.
Addressing these two dimensions (unrestricted training data and strict temporal inference separation) is non-trivial and often overlooked in current evaluation standards. We ask the community to develop ideas for new benchmark methodologies which are grounded in these two requirements. Possible directions could be constantly newly sourced test data, synthetic test data, or fast-paced competitions on unreleased and unbiased test data, just to name a few.
5 Conclusion
Remarkably, TSFMs appear to follow a development trajectory similar to that of LLMs in terms of both data and model size. Scaling laws seem to apply to TSFMs, and there are notable similarities in the evaluation challenges they face, particularly concerning potential information leakage. We highlight key evaluation challenges that arise from the emergence of TSFMs. By investigating the data lineage of 22 prominent TSFMs and reflecting on common TSFM training and evaluation strategies, we identified two significant sources of information leakage: direct information leakage by train-test sample overlaps through multi-purpose use of datasets and indirect information leakage by temporal overlap of correlated train and test series.
Current benchmark approaches from time series forecasting or machine learning do not mitigate all information leakage risks. Based on our investigations and existing methodologies, we propose two key requirements for information leakage free evaluations in the age of TSFMs: flexibility in pre-training data and strict global post-training test periods. Beyond the development of new benchmark methodologies, we identify several directions for future research . To date, the effect size of test-set contamination—when a test dataset is present in the pre-training corpus—has only been shown to be significant in isolated cases. Likewise, the magnitude of temporal leakage arising from correlated and overlapping time series remains largely unknown in real TSFM settings and is likely to vary across domains, types of global events, and other contextual factors. Carefully designed TSFM training regimes will be required to rigorously quantify these effects.
Ultimately, we believe that the requirements we have established lay the foundation for robust, fair, and information-leakage-free benchmarking. This is essential for preventing an evaluation crisis for TSFMs akin to what has been observed with LLMs.
6 Data Availability
The resulting table of the TSFM data-lineage analysis is supplied in the Supplementary Information Section D. For the global pattern experiment in Section 3.2, the used data is available through the source code via zenodo: https://doi.org/10.5281/zenodo.18635478
7 Code Availability
The code to reproduce the experiment in Section 3.2 is provided via zenodo: https://doi.org/10.5281/zenodo.18635478
Appendix A Time Series Foundation Model Analysis
A.1 Multi-Purpose Use of Train and Test Data
This section showcases the varying training and testing dataset strategies of selected TSFMs. Table 3 reveals that there is virtually no standardized approach among TSFMs, with each model employing different combinations of datasets for train/test (or in-domain) datasets. While some datasets like Electricity (H) and Traffic Hourly (H) appear across multiple models, the coverage remains fragmented. Table 4 further illustrates this inconsistency in zero-shot evaluation practices. Even among datasets with five or more overlaps, the specific models testing on each dataset vary considerably. For instance, ETT2 (15T) is evaluated by Chronos, Moirai, and Lag-Llama, while Electricity (H) appears in Moirai, Moirai-MoE, and LLMTime.
Another notable case for the multi-purpose use of datasets is the subset Australian Electricity Demand from the Monash dataset collection (Godahewa et al., 2021), which is used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM). This multi-purpose use makes the dataset practically useless for benchmarking between models (e.g., it cannot be used to compare Chronos to Lag-Llama). To understand the problem of overlaps better, we provide Figure 6.
Also an interesting case is the Informer collection (Zhou et al., 2021), which is primarily used for zero-shot evaluations. However, some individual time series of the Informer collection, such as ETTh1, ETTh2, and ETTm1, have also been used for pre-training (e.g., in LagLlama) or train/test evaluation (e.g., in UniTime). This means that LagLlama or UniTime cannot be benchmarked against any other model using the Informer collection.
Taken together with investigations of Roque et al. (2025) that the selection of evaluation datasets can highly influence the final evaluation results, this strengthens the need for a more uniform benchmark approach.
Figure 6: Test set contaminations when comparing TSFM A with TSFM B with overlapping pre-training and test sets on the test set of TSFM A.
Table 3: Comparison of train/test evaluations of selected different models ordered by the total number of overlaps. For full results see Section D.
Dataset Chronos Moirai MOMENT Time-LMM
Electricity (H) T/T - T/T T/T
Traffic Hourly (H) - T/T T/T T/T
ETT1 (H) - - T/T T/T
Weather (Australia) (D) - T/T - T/T
M4 Monthly (M) T/T T/T - -
M4 Hourly (H) T/T T/T - -
M4 Daily (D) T/T T/T - -
M4 Weekly (W-SUN) T/T T/T - -
ETT1 (15T) - - T/T T/T
ETT2 (15T) - - T/T T/T
ETT2 (H) - - T/T T/T
CDC Fluview ILINet (W) - - T/T T/T
Table 4: Comparison of zero-shot evaluations of selected different models ordered by the total number of overlaps. Only datasets with 5 or more overlaps are shown. Models which have not included any of the shown datasets into their train/test evaluation are not displayed. For full results see Section D.
Dataset Chronos Moirai Moirai-MoE Lag-Llama LLMTime
Jena Weather (10T) - ZS ZS - ZS
ETT2 (15T) ZS ZS - ZS -
Electricity (H) - ZS ZS - ZS
ETT2 (H) ZS ZS - - -
ETT1 (H) ZS ZS - - -
ETT1 (15T) ZS ZS - - -
Traffic Hourly (H) ZS - - - ZS
Appendix B Leakage Investigations
B.1 Effects of Overlaps in Train-Test Samples
This section collects different examples where the train-test sample overlaps unintentionally led to information leakage in recent benchmark papers. In one case, the benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used in the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b). Based on our analysis of the benchmark results, this leads to an advantage of 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets. The advantage of the best TSFM on non-leaked datasets is only between 0.3% and 14%. The performance benefit through information leakage appears to align with findings from the Moirai leakage example (see Section B.2). In another example a peer-reviewed paper benchmarked TSFM models on the Electricity dataset which is included in every tested TSFM pretraining data Saravanan et al. (2024). Also in the energy domain, a paper benchmarked TSFMs on the Spanish dataset (among others), which is also included in the pretraining data of the tested TSFMs. Chronos achieved a sMAPE of 4.854 where the best non-TSFM model TiDE achieved a sMAPE of 8.102 Montet et al. (2025). This highlights the increasing difficulties for benchmarking TSFMs.
B.2 Intended Information-Leakage Experiment for Moirai Model
Table 5: Information leakage in TSFM comparing Moirai with (MoiLeak) and without (Moi) information leakage during training phase for different model sizes. Metric is MAPE. Based on experiments by Aksu et al. (2024).
Horizon Short Medium Long
Moirai Leakage Moirai Moirai Leakage Moirai Moirai Leakage Moirai
Dataset Frequency Model Size
loop_seattle 5T S 0.84 0.87 0.75 0.77 0.70 0.75
B 0.67 0.84 0.42 0.83 0.50 0.78
L 0.66 0.83 0.33 0.85 0.46 0.81
loop_seattle H S 1.22 1.19 0.73 0.70 0.70 0.71
B 0.96 1.08 0.54 0.65 0.49 0.59
L 0.84 0.89 0.53 0.71 0.47 1.18
m_dense H S 0.73 0.70 0.70 0.71 0.71 0.91
B 0.54 0.65 0.49 0.59 0.53 0.61
L 0.53 0.71 0.47 1.18 0.49 1.69
sz_taxi 15T S 0.95 1.11 0.65 0.60 2.12 2.30
B 0.90 0.84 0.71 0.64 2.16 2.42
L 0.78 0.82 0.71 0.60 2.14 2.24
The empirical analysis using the Moirai TSFM by Aksu et al. (2024) demonstrates the effects of information leakage between training and test sets on the predictive performance of TSFMs. The authors prepared a new pre-training and held-out evaluation dataset without any overlap and trained the Moirai architecture on it. A different Moirai model from a previous publication was taken which pre-training dataset contained 0.1% data of the newly defined held-out evaluation dataset leading to a deliberate small portion of information-leakage. The investigated data includes nine datasets tested on three different forecast horizons (short, medium, long). Not for all horizons data was available, so we filtered the results on 4 datasets where the results are available for all horizons to reasonably compare the metrics and model sizes, as shown in Figure 7. On short horizons, both models, leakage-free and leakage model, perform better with increasing model size while Moirai Leakage has on average 7.6% points lower MAPE score. At medium horizons, the leakage-free MoiraiL generally underperforms compared to other model sizes, exhibiting an average performance deficit of 32% points compared to the Moirai LeakageL model. On long horizons, all models perform significantly worse. However, Moirai Leakage achieves an average MAPE score that is 29% points lower than the leakage-free Moirai. Across all horizons, the biggest difference is between MoiraiL and Moirai LeakageL.
These results vividly demonstrate the significant impact that train-test sample overlap has on the performance of TSFMs. Larger model sizes especially seem prone to this.
Figure 7: Comparing the model Moirai once with information leakage (Moirai Leakage), where the model was exposed to the test data during pre-training and the same model trained without information leakage (Moirai). Short, Medium and Long forecasting horizons were investigated. Model sizes are S (small), B (base) and L (large) and compared by the MAPE metric (the lower the better). Based on experiments by Aksu et al. (2024).
B.3 Potential Overlaps of Correlated Train and Test Series in Dataset Collections
It may be theoretically possible that existing model evaluation results have already been influenced by overlaps of correlated train and test series. We analyse the overlapping time frames in the Monash Dataset  [^10] https://huggingface.co/datasets/Monash-University/monash_tsf as can be seen in Figure 8. Test datasets shouldn’t lie in the same time period as other training datasets. As an example, the proposed Australian Electricity Demand test time frame (April 2015) lays in the same period as the training period of the "Weather (Australia)" training set (1900-2021). The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if their are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage, at least on theoretical basis. Similarly, the Oikolab Weather dataset ranges from 2010 January to 2021 May, which is weather data located in Melbourne. The test set of the dataset Pedestrian Count (also Melbourne) is located at the end of April 2020. This test period lays within the training data of the Oikolab Weather, which could be a greater potential global pattern information leakage than in previous example. Both datasets are included in the datasets of Moirai, Time-Moe, Sundial and Toto.
Moreover, LagLlama uses the Beijing Multi-Site Air Quality dataset from 12 districts in Beijing from 2013 to 2017 as a pre-training dataset and simultaneously uses the Beijing PM2.5 air quality dataset from the Beijing airport between 2010 and 2013, which could also in theory lead to leakage.
Figure 8: Date range of datasets within Monash dataset collection. Note that the “cif_2016" dataset was originally delivered without a timestamp and the start year was defined as 1900. No data cleaning was done.
Appendix C Temporal Overlap of Correlated Series Experiment
This section provides implementation details of the experiment conducted to show the influence of the temporal overlap of correlated train and test series in a TSFM setting.
C.1 Data visualizations and statistics
Figure 9 shows two example time series from major stock indices, the DAX and S&P 500. Despite representing different markets, both series display remarkably similar patterns during key economic phases, including crisis periods and growth cycles. This visual comparison highlights how global financial markets often move in tandem, reflecting shared responses to worldwide economic events.
Figure 10 illustrates the relationship between training data and test data that overlap temporally. It is important to note that while the training and test periods overlap in time, they do not share the same data points, the overlap is purely temporal, meaning both datasets cover in a small part the same time period but contain different observations or series as can be seen in Table 6.
Figure 9: Comparison of DAX and S&P 500 points over time
Figure 10: Comparison of time periods per dataset
Category No Leakage Indirect Leakage
Train Test Train Test
# Time Series 484 1 484 1
# Data Points 205,854 230 205,859 230
# Indirect Leakage Data Points - - 1,600 -
Table 6: Summary of time series and data points with and without leakage
C.2 Scaling and Model Tuning
Min-max scaling was applied separately to each dataset to account for differences in scale across datasets and ensure optimal model performance. Each scaler was fitted on the respective training data and subsequently used to transform the remaining data. For the S&P 500 dataset, the scaler was fitted explicitly on data from 2015 to the end of 2019 and then applied to the data from 2020. The Time-Series Transformer was implemented via the tsai library [^11] https://timeseriesai.github.io/tsai. The tuning was performed via Optuna [^12] https://github.com/optuna/optuna for 50 trials to minimize the MAE on the validation set. The training set consisted of the first 80% of NN5 Daily and Tourism Monthly datasets as well as January 1st, 2015 until June 30th, 2019 for no-leak and January 1st 2016 until June 30th, 2020 of the stock index data for leak. The validation set consisted of the remaining 20% of NN5 Daily and Tourism Monthly datasets, and the stock index data from July 1st, 2019 until December 31st, 2019 for no-leak and July 1st, 2020 until December 31st 2020 for leak.
Parameter Lower Upper Stepsize Choice
Learning rate log(1e-05) log(1e-01) - -
Batch size - - - 16, 32, 64, 128
Epochs 5 30 5 -
Dropout 0.0 0.7 - -
FC Dropout 0.0 0.7 - -
Number of layers 2 8 1 -
Model dimension - - - 64, 128, 256, 512, 1024
Number of heads - - - 4, 8, 16, 32
Dimension of feedforward network model - - - 32, 64, 128, 256, 512
Table 7: Optuna search space for tuning
Parameter No-leak Leak
Learning rate 1.018e-03 1.431e-03
Batch size 128 128
Epochs 30 30
Dropout 0.5 0.3
FC Dropout 0.27 0.004
Number of layers 2 4
Model dimension 128 128
Number of heads 32 4
Dimension of feedforward network model 32 512
Table 8: Final model parameters
Appendix D TSFM Data Lineage
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
Jena Weather Autoformer (Wu et al., 2021) Nature 10T ZS ZS T/T T/T ZS ZS ZS ZS T/T ZS T/T ZS T/T ZS T/T
Jena Weather Autoformer (Wu et al., 2021) Nature H ZS T/T ZS
Jena Weather Autoformer (Wu et al., 2021) Nature D ZS ZS T/T ZS
BizITObs - Application AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS ZS T/T ZS
BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps 5T ZS ZS T/T ZS
BizITObs - L2C AutoMixer (Palaskar et al., 2024) Web/CloudOps H ZS T/T ZS
BizITObs - Service AutoMixer (Palaskar et al., 2024) Web/CloudOps 10S ZS T/T ZS
BDG-2 Bear BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
BDG-2 Fox BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
BDG-2 Panther BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
BDG-2 Rat BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
Borealis BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
Buildings900K BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
IDEAL BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
Low Carbon London BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P
Sceaux BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
SMART BuildingsBench (Emami et al., 2023) Energy H P P P P P P P P P P
CMIP6 ClimateLearn (Nguyen et al., 2023) Climate 6H P P P P P P P P
ERA5 ClimateLearn (Nguyen et al., 2023) Climate H P P P P P P P P P
Alibaba Cluster Trace 2018 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P
Azure vM Traces 2017 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P P P P
Borg Cluster Data 2011 CloudOpsTSF (Woo et al., 2023) CloudOps 5T P P P P P P P P P P
Air Passengers Darts Dataset (Herzen et al., 2021) Econ/Fin M ZS ZS
Aus Beer Darts Dataset (Herzen et al., 2021) Econ/Fin Q ZS ZS
Gas Rate CO2 Darts Dataset (Herzen et al., 2021) Econ/Fin Y ZS ZS
Monthly Milk Darts Dataset (Herzen et al., 2021) Econ/Fin M ZS ZS
Wine Darts Dataset (Herzen et al., 2021) Sales M ZS ZS
Wooly Darts Dataset (Herzen et al., 2021) Econ/Fin Q ZS ZS
Heart Rate Darts Dataset (Herzen et al., 2021) Nature 0.5S ZS ZS
Sunspots Darts Dataset (Herzen et al., 2021) Nature M ZS
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T ZS T/T ZS
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H ZS T/T ZS
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps 5T ZS T/T ZS
Bitbrains Fast Storage Grid Workloads Archive (Shen et al., 2015) Web/CloudOps H ZS T/T ZS
FRED FRED (Oreshkin et al., 2021) Econ/Fin D P
FRED FRED (Oreshkin et al., 2021) Econ/Fin Y P
FRED FRED (Oreshkin et al., 2021) Econ/Fin W P
FRED FRED (Oreshkin et al., 2021) Econ/Fin M P
FRED FRED (Oreshkin et al., 2021) Econ/Fin Q P
M5 GluonTS (Alexandrov et al., 2020a) Sales D ZS P P P P P P P ZS ZS
Taxi GluonTS (Alexandrov et al., 2020a) Transport 30T T/T P P P P P T/T P P P P
Taxi GluonTS (Alexandrov et al., 2020a) Transport H P P possible P P P P P P
Uber TLC Daily GluonTS (Alexandrov et al., 2020a) Transport D T/T P P P P P T/T P P P
Uber TLC Hourly GluonTS (Alexandrov et al., 2020a) Transport H T/T P P P P P P T/T P P P
Wiki-Rolling GluonTS (Alexandrov et al., 2020a) Web D P P P P
Trends hourly Google Trends Web H P P
Trends weekly Google Trends Web W P P
Trends monthly Google Trends Web D P P
Requests Huawei Cloud (Joosen et al., 2023) CloudOps T ZS
Function Delay Huawei Cloud (Joosen et al., 2023) CloudOps T P
Platform Delay Huawei Cloud (Joosen et al., 2023) CloudOps T ZS
CPU Usage Huawei Cloud (Joosen et al., 2023) CloudOps T P
Memory Usage Huawei Cloud (Joosen et al., 2023) CloudOps T P
CPU Limit Huawei Cloud (Joosen et al., 2023) CloudOps T P
Memory Limit Huawei Cloud (Joosen et al., 2023) CloudOps T P
Instances Huawei Cloud (Joosen et al., 2023) CloudOps T P
ETT1 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS ZS T/T
ETT1 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS ZS T/T
ETT1 Informer (Zhou et al., 2020) Energy D ZS ZS ZS T/T ZS
ETT1 Informer (Zhou et al., 2020) Energy W-THU ZS T/T ZS
ETT2 Informer (Zhou et al., 2020) Energy 15T ZS ZS T/T ZS T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS T/T
ETT2 Informer (Zhou et al., 2020) Energy H ZS ZS T/T P T/T ZS ZS ZS T/T T/T T/T ZS T/T ZS T/T ZS ZS T/T
ETT2 Informer (Zhou et al., 2020) Energy D ZS ZS ZS T/T ZS
ETT2 Informer (Zhou et al., 2020) Energy W-THU ZS T/T ZS
Weather (US) Informer (Zhou et al., 2020) Weather H ZS T/T
Brazilian Cities Temperature Kaggle Competition Weather M P possible P P P P P
Walmart Store Sales Kaggle Competition Sales W ZS ZS ZS
Istanbul Traffic Kaggle Competition Transport H ZS ZS ZS
Turkey Power Kaggle Competition Energy H ZS ZS ZS ZS
TSMC Stock Kaggle Competition Econ/Fin D ZS
LargeST LargeST (Liu et al., 2023a) Transport 5T P P P P P P P P P P
Beijing Subway LibCity (Wang et al., 2023a) Transport 30T P P P P P P P P P
HZMetro LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P
Loop Seattle LibCity (Wang et al., 2023a) Transport 5T P P P ZS P T/T ZS
Loop Seattle LibCity (Wang et al., 2023a) Transport H ZS T/T ZS
Loop Seattle LibCity (Wang et al., 2023a) Transport D ZS T/T ZS
Los-Loop LibCity (Wang et al., 2023a) Transport 5T P P P P P P P T/T ZS P
M_DENSE LibCity (Wang et al., 2023a) Transport H ZS T/T ZS
M_DENSE LibCity (Wang et al., 2023a) Transport D ZS ZS T/T ZS
PEMS Bay LibCity (Wang et al., 2023a) Transport 5T P P P P P P P P P P P P
PEMS03 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
PEMS04 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
PEMS07 LibCity (Wang et al., 2023a) Transport 5T P P ZS P P P P P P P P P P
PEMS08 LibCity (Wang et al., 2023a) Transport 5T P P T/T ZS P P P P P P P P P P
Q-Traffic LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P P P P
SHMetro LibCity (Wang et al., 2023a) Transport 15T P P P P P P P P P P P
SZ-Taxi LibCity (Wang et al., 2023a) Transport 15T P P P ZS T/T ZS P
SZ-Taxi LibCity (Wang et al., 2023a) Transport H P ZS T/T ZS
Rotterdam LibCity (Wang et al., 2023a) Transport 2T P P P T/T ZS
Beijing Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P P P P P P P P P P
CDC Fluview ILINet LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T ZS T/T T/T P P P P P P P P
CDC Fluview WHO NREVSS LOTSA_Others (Woo et al., 2024) Healthcare W P P ZS T/T T/T P P P P P P P P
China Air Quality LOTSA_Others (Woo et al., 2024) Nature H P P P P P P P P P P P
Favorita Sales LOTSA_Others (Woo et al., 2024) Sales D P P P P P P P P P P P
Favorita Transactions LOTSA_Others (Woo et al., 2024) Sales D P P P P P P P
GoDaddy LOTSA_Others (Woo et al., 2024) Econ/Fin M P P P P P P P P P
KDD Cup 2022 LOTSA_Others (Woo et al., 2024) Energy 10T P P P P P P P P P P
Project Tycho LOTSA_Others (Woo et al., 2024) Healthcare W P P P P P P P
Residential Load Power LOTSA_Others (Woo et al., 2024) Energy T P P P P P P P P P P
Residential PV Power LOTSA_Others (Woo et al., 2024) Energy T P P P P P P P P P P
Restaurant LOTSA_Others (Woo et al., 2024) Sales D P
Solar LSTNet (Lai et al., 2017) Energy 10T ZS T/T T/T ZS T/T
Solar LSTNet (Lai et al., 2017) Energy 5T P possible P P P P
Solar LSTNet (Lai et al., 2017) Energy H P ZS ZS ZS ZS P P T/T ZS P
Solar LSTNet (Lai et al., 2017) Energy D ZS T/T ZS
Solar LSTNet (Lai et al., 2017) Energy W-FRI ZS T/T ZS
Exchange Rate LSTNet (Lai et al., 2017) Econ/Fin D ZS ZS T/T ZS ZS T/T P T/T ZS ZS
Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales D P ZS T/T ZS
Hierarchical Sales (M5 Competition) Mancuso et al. (2020) Sales W-WED ZS T/T ZS
Australian Electricity Demand Monash (Godahewa et al., 2021) Energy 30T ZS ZS T/T T/T P P P ZS P ZS P P P T/T P P ZS ZS P
Bitcoin Monash (Godahewa et al., 2021) Econ/Fin D ZS T/T T/T P P ZS P ZS P P P T/T P P P
Car Parts Monash (Godahewa et al., 2021) Sales M T/T T/T P ZS P ZS T/T T/T ZS ZS ZS
CIF 2016 Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P P T/T P P ZS ZS
COVID Deaths Monash (Godahewa et al., 2021) Healthcare D ZS ZS T/T T/T P ZS ZS P ZS T/T T/T ZS ZS ZS
Covid Mobility Monash (Godahewa et al., 2021) Transport D T/T T/T P P T/T P P
Elecdemand Monash (Godahewa et al., 2021) Energy 30T T/T T/T P P T/T P P
Extended Web Traffic Monash (Godahewa et al., 2021) Web D T/T T/T P P P P T/T P P
FRED MD Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS T/T T/T P ZS ZS P P P T/T P P ZS ZS P
Hospital Monash (Godahewa et al., 2021) Healthcare M ZS ZS T/T T/T P ZS ZS P ZS T/T T/T ZS ZS ZS
Kaggle Web Traffic Weekly Monash (Godahewa et al., 2021) Web W P P P P P P P P P P P
KDD Cup 2018 Monash (Godahewa et al., 2021) Nature H T/T T/T T/T P P P ZS P P ZS T/T T/T T/T ZS P
KDD Cup 2018 Monash (Godahewa et al., 2021) Nature D ZS P T/T ZS P
London Smart Meters Monash (Godahewa et al., 2021) Energy 30T T/T P P P P P P P P P P T/T P P P P P
M1 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T P ZS P P T/T P P ZS ZS
M1 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P P P P P P P ZS ZS
M1 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P P P P P P P ZS ZS
M3 Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS T/T T/T ZS T/T ZS P P T/T P P ZS ZS
M3 Other Monash (Godahewa et al., 2021) Econ/Fin Q T/T T/T T/T ZS P T/T P P
M3 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS P P ZS T/T P P P P P ZS ZS
M3 Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS P P ZS T/T P P P P P ZS ZS
M4 Daily Monash (Godahewa et al., 2021) Econ/Fin D T/T P T/T T/T P T/T ZS P ZS T/T T/T T/T ZS P P
M4 Hourly Monash (Godahewa et al., 2021) Econ/Fin H T/T P T/T T/T P T/T ZS P ZS T/T T/T T/T ZS P P
M4 Monthly Monash (Godahewa et al., 2021) Econ/Fin M T/T P T/T T/T ZS T/T ZS P ZS T/T T/T T/T ZS P P
M4 Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q-DEC ZS P P P ZS T/T P ZS P T/T ZS ZS P
M4 Weekly Monash (Godahewa et al., 2021) Econ/Fin W-SUN T/T P T/T T/T T/T ZS P ZS T/T T/T T/T ZS P
M4 Yearly Monash (Godahewa et al., 2021) Econ/Fin A-DEC ZS P P P ZS T/T P ZS P T/T ZS ZS P
NN5 Daily Monash (Godahewa et al., 2021) Econ/Fin D ZS ZS T/T T/T P ZS P ZS P P P T/T P P ZS ZS P
NN5 Weekly Monash (Godahewa et al., 2021) Econ/Fin W ZS ZS T/T T/T P ZS ZS P T/T P P ZS ZS
Oikolab Weather Monash (Godahewa et al., 2021) Climate H T/T T/T P P T/T P P
Pedestrian Counts Monash (Godahewa et al., 2021) Transport H T/T ZS T/T T/T P ZS P ZS ZS P P P T/T T/T P P P P
Rideshare Monash (Godahewa et al., 2021) Transport H T/T T/T T/T P ZS P P T/T T/T P P P
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
Saugeen Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P ZS P ZS P ZS P T/T T/T ZS P
Saugeen Monash (Godahewa et al., 2021) Nature W-THU ZS T/T ZS
Saugeen Monash (Godahewa et al., 2021) Nature M ZS T/T ZS
Solar Power (Australia) Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P P P T/T P P P P P
Solar Monash (Godahewa et al., 2021) Energy W ZS P P ZS ZS
Solar Monash (Godahewa et al., 2021) Energy 10M P P P ZS
Sunspot Monash (Godahewa et al., 2021) Nature D ZS T/T T/T P P P ZS P ZS P P P T/T P P P
Temperature Rain Monash (Godahewa et al., 2021) Nature D T/T T/T T/T P P ZS P ZS P T/T T/T T/T ZS P P
Tourism Monthly Monash (Godahewa et al., 2021) Econ/Fin M ZS ZS P T/T ZS ZS P P P P ZS ZS
Tourism Quarterly Monash (Godahewa et al., 2021) Econ/Fin Q ZS ZS P T/T ZS ZS P P P P ZS ZS
Tourism Yearly Monash (Godahewa et al., 2021) Econ/Fin Y ZS ZS P P P T/T ZS P P P P P ZS ZS
Traffic Hourly Monash (Godahewa et al., 2021) Transport H ZS P T/T T/T T/T P T/T ZS ZS ZS ZS T/T P P T/T T/T P P ZS P T/T
Traffic Weekly Monash (Godahewa et al., 2021) Transport W ZS T/T T/T ZS P ZS ZS P P T/T P P
US Births Monash (Godahewa et al., 2021) Healthcare D ZS T/T T/T P P ZS P ZS P ZS P T/T T/T P
US Births Monash (Godahewa et al., 2021) Healthcare W-TUE ZS T/T ZS
US Births Monash (Godahewa et al., 2021) Healthcare M ZS T/T ZS
Vehicle Trips Monash (Godahewa et al., 2021) Transport D T/T T/T P ZS P P T/T P P
Weather (Australia) Monash (Godahewa et al., 2021) Climate D ZS ZS T/T T/T P ZS ZS P T/T T/T P P P T/T P P ZS P
Wind Farms Monash (Godahewa et al., 2021) Energy T T/T T/T P P P P T/T P P P
Wind Farms Monash (Godahewa et al., 2021) Energy H P P P P P
Wind Farms Monash (Godahewa et al., 2021) Energy D P possible P P P P
Wind Farms Monash (Godahewa et al., 2021) Energy 5T P possible P
Wind Power Monash (Godahewa et al., 2021) Energy 4S T/T T/T P P P P T/T P P P P P P
Kaggle Web Traffic Daily Monash (Godahewa et al., 2021) Web D P P P P
Restaurant Recruit Other Sales D ZS T/T ZS
Subseasonal Other Climate D P P P P P P P P P P
Subseasonal Precipitation Other Climate D P P P P P P P P P
USHCN Other Weather D P possible P P P P P
Mexico City Bikes Other Transport H P P possible P P P P P
Dominick Other Sales D ZS P P P ZS ZS
ERCOT Load Other Energy H ZS P P ZS ZS
Weather (US) Other Weather 10T P ZS
Bike Sharing Other Transport H ZS ZS
CC Carbon Other Energy 2T ZS
Air Quality Other Nature H P
Beijing PM 2.5 Other Nature H ZS
TDBrain Other Healthcare 0.002S P P P
SensorData Other IoT 0.02S P
BDG-2 Bull ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
BDG-2 Cockatoo ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
BDG-2 Hog ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
Covid19 Energy ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
ELF ProEnFo (Wang et al., 2023b) Energy H P P P P P P
GEF12 ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
GEF14 ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
GEF17 ProEnFo (Wang et al., 2023b) Energy H P P P P P P
PDB ProEnFo (Wang et al., 2023b) Energy H P P P P P P P
Spanish ProEnFo (Wang et al., 2023b) Energy H P P P P P P P P P P
Australia Rainfall Tan et al., 2021 Nature H P
Beijing PM 2.5 Tan et al., 2021 Nature H P
BenzeneConcentration Tan et al., 2021 Nature H P
IEEEPPG Tan et al., 2021 Healthcare 0.008S P
BIDMC32HR Tan et al., 2021 Healthcare - P
Dodger TSB-UAD (Paparrizos et al., 2022) Traffic - P
ECG TSB-UAD (Paparrizos et al., 2022) Healthcare - P P
IOPS TSB-UAD (Paparrizos et al., 2022) Web - P
KDD21 TSB-UAD (Paparrizos et al., 2022) Multiple - P
MGAB TSB-UAD (Paparrizos et al., 2022) Synthetic - P
NAB TSB-UAD (Paparrizos et al., 2022) Web - P
SensorScope TSB-UAD (Paparrizos et al., 2022) Nature - P
YAHOO TSB-UAD (Paparrizos et al., 2022) Web - P
NASA-MSL TSB-UAD (Paparrizos et al., 2022) Sensor - P
NASA-SMAP TSB-UAD (Paparrizos et al., 2022) Sensor - P
Daphnet TSB-UAD (Paparrizos et al., 2022) Gait - P
GHL TSB-UAD (Paparrizos et al., 2022) Sensor - P
Genesis TSB-UAD (Paparrizos et al., 2022) Sensor - P
MITDB TSB-UAD (Paparrizos et al., 2022) Healthcare - P
OPP TSB-UAD (Paparrizos et al., 2022) Sensor - P
Occupancy TSB-UAD (Paparrizos et al., 2022) Sensor - P
SMD TSB-UAD (Paparrizos et al., 2022) Web - P
SZSDB TSB-UAD (Paparrizos et al., 2022) Healthcare - P
Electricity UCI ML Archive (Trindade, 2015) Energy 15T T/T P T/T P ZS T/T T/T T/T ZS P
Electricity UCI ML Archive (Trindade, 2015) Energy H T/T P ZS ZS T/T P T/T ZS ZS ZS ZS T/T T/T ZS T/T ZS T/T T/T ZS P P T/T
Electricity UCI ML Archive (Trindade, 2015) Energy D ZS ZS T/T ZS
Electricity UCI ML Archive (Trindade, 2015) Energy W-FRI T/T ZS T/T T/T ZS
Electricity UCI ML Archive (Trindade, 2015) Energy M T/T T/T
Electricity UCI ML Archive (Trindade, 2015) Energy W P P P
SemgHandGenderCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
GestureMidAirD2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
UWaveGestureLibraryAll UCR Archive (Ismail Fawaz et al., 2019) Human - T/T
SelfRegulationSCP1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
UWaveGestureLibraryX UCR Archive (Ismail Fawaz et al., 2019) Human - P
GesturePebbleZ2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
Healthcare5000 (orignally: ECG5000) UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T P
OSULeaf UCR Archive (Ismail Fawaz et al., 2019) Nature - T/T
MedicalImages UCR Archive (Ismail Fawaz et al., 2019) Healthcare - T/T
Haptics UCR Archive (Ismail Fawaz et al., 2019) Human - P
LargeKitchenAppliances UCR Archive (Ismail Fawaz et al., 2019) Power - P
Japanesevowels UCR Archive (Ismail Fawaz et al., 2019) Audio - P
Worms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P
Ham UCR Archive (Ismail Fawaz et al., 2019) Facilities - T/T
DistalPhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ProximalPhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
SemgHandMovementCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
RefrigerationDevices UCR Archive (Ismail Fawaz et al., 2019) Power - P
FreezerRegularTrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
PigAirwayPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P
TwoLeadECG UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
GunPointMaleversusFemale UCR Archive (Ismail Fawaz et al., 2019) Human - P
Trace UCR Archive (Ismail Fawaz et al., 2019) Power - P
SmoothSubspace UCR Archive (Ismail Fawaz et al., 2019) Generated - P
MiddlePhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
AtrialFibrillation UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
SyntheticControl UCR Archive (Ismail Fawaz et al., 2019) Generated - P
ShapesAll UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Human UCR Archive (Ismail Fawaz et al., 2019) BodyZSerticalSignal - P
PLAID UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
AllGestureWiimoteX UCR Archive (Ismail Fawaz et al., 2019) Human - P
Heartbeat UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
Wafer UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
FaceFour UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Phoneme UCR Archive (Ismail Fawaz et al., 2019) Audio - P P P P
InlineSkate UCR Archive (Ismail Fawaz et al., 2019) Human - P
CricketX UCR Archive (Ismail Fawaz et al., 2019) Human Body - P
SelfRegulationSCP2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
DistalPhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ChlorineConcentration UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Chinatown UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
GestureMidAirD1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
MiddlePhalanxOutlineAgeGroup UCR Archive (Ismail Fawaz et al., 2019) Nature - P
UMD UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Crop UCR Archive (Ismail Fawaz et al., 2019) Nature - P
PenDigits UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
GesturePebbleZ1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
Handwriting UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
Mallat UCR Archive (Ismail Fawaz et al., 2019) Generated - P
ERing UCR Archive (Ismail Fawaz et al., 2019) Human - P
StarLightCurves UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
WordSynonyms UCR Archive (Ismail Fawaz et al., 2019) Audio - P
PEMS-SF UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
FaceDetection UCR Archive (Ismail Fawaz et al., 2019) Human - P
Computers UCR Archive (Ismail Fawaz et al., 2019) Power - P
ArrowHead UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Wine UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Coffee UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Earthquakes UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Herring UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Lightning2 UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Beef UCR Archive (Ismail Fawaz et al., 2019) Nature - P
MiddlePhalanxOutlineCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
HealthcareFiveDays UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
Yoga UCR Archive (Ismail Fawaz et al., 2019) Human - P
Adiac UCR Archive (Ismail Fawaz et al., 2019) Nature - P
HandMovementDirection UCR Archive (Ismail Fawaz et al., 2019) Human - P
Dataset Source Domain Frequency Chronos TimesFM Moirai Moirai-MoE Time-GPT ForecastPFN MOMENT Lag-Llama GPT4TS Timer VisionTS TTM LLMTime Time-LMM UniTime Time-MoE Chronos-2 Flowstate(under review) LightGTS Sundial TabPFN-TS Time-MoE2 TinyTimeMixer TiRex Toto YingLong (under review) Gift-Eval TOTO2 Flowstate(T) TimesFM-ICL ROSE
MoteStrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
Rock UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Strawberry UCR Archive (Ismail Fawaz et al., 2019) Nature - P
InsectWingbeatSound UCR Archive (Ismail Fawaz et al., 2019) Nature - P
DodgerLoopWeekend UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
MixedShapesSmallTrain UCR Archive (Ismail Fawaz et al., 2019) Generated - P
EthanolConcentration UCR Archive (Ismail Fawaz et al., 2019) Power - P
OliveOil UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Meat UCR Archive (Ismail Fawaz et al., 2019) Nature - P
MelbournePedestrian UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
Car UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
FaceAll UCR Archive (Ismail Fawaz et al., 2019) Human - P
FacesUCR UCR Archive (Ismail Fawaz et al., 2019) Human - P
AllGestureWiimoteY UCR Archive (Ismail Fawaz et al., 2019) Human - P
NATOPS UCR Archive (Ismail Fawaz et al., 2019) Human - P
SemgHandSubjectCh2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
ShakeGestureWiimoteZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
Cricket UCR Archive (Ismail Fawaz et al., 2019) Human - P
BME UCR Archive (Ismail Fawaz et al., 2019) Generated - P
EigenWorms UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
FordB UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
NonInvasiveFetalHealthcareThorax1 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
UWaveGestureLibrary UCR Archive (Ismail Fawaz et al., 2019) Human - P
CinCHealthcareTorso UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
PigArtPressure UCR Archive (Ismail Fawaz et al., 2019) Nature - P P P P
Fish UCR Archive (Ismail Fawaz et al., 2019) Nature - P
SonyAIBORobotSurface2 UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
FiftyWords UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
MotorImagery UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
ToeSegmentation1 UCR Archive (Ismail Fawaz et al., 2019) Human - P
PhonemeSpectra UCR Archive (Ismail Fawaz et al., 2019) Audio - P
FreezerSmallTrain UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
TwoPatterns UCR Archive (Ismail Fawaz et al., 2019) Generated - P
ShapeletSim UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Plane UCR Archive (Ismail Fawaz et al., 2019) Generated - P
GestureMidAirD3 UCR Archive (Ismail Fawaz et al., 2019) Human - P
DiatomSizeReduction UCR Archive (Ismail Fawaz et al., 2019) Generated - P
Human UCR Archive (Ismail Fawaz et al., 2019) BodyHorizontalSignal - P
CricketZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
StandWalkJump UCR Archive (Ismail Fawaz et al., 2019) Human - P
WormsTwoClass UCR Archive (Ismail Fawaz et al., 2019) Human - P
Lightning7 UCR Archive (Ismail Fawaz et al., 2019) Nature - P
UWaveGestureLibraryY UCR Archive (Ismail Fawaz et al., 2019) Human - P
GunPointAgeSpan UCR Archive (Ismail Fawaz et al., 2019) Human - P
DistalPhalanxOutlineAgeGroup UCR Archive (Ismail Fawaz et al., 2019) Nature - P
SwedishLeaf UCR Archive (Ismail Fawaz et al., 2019) Nature - P
LSST UCR Archive (Ismail Fawaz et al., 2019) Nature - P
CBF UCR Archive (Ismail Fawaz et al., 2019) Generated - P
BeetleFly UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Libras UCR Archive (Ismail Fawaz et al., 2019) Human - P
HouseTwenty UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
ScreenType UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
InsectEPGSmallTrain UCR Archive (Ismail Fawaz et al., 2019) Nature - P
AllGestureWiimoteZ UCR Archive (Ismail Fawaz et al., 2019) Sensor - P
DodgerLoopDay UCR Archive (Ismail Fawaz et al., 2019) Sensor - P
NonInZSasiZSeFetalHealthcareThorax2 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
BasicHuman UCR Archive (Ismail Fawaz et al., 2019) Bodys - P
GunPointOldversusYoung UCR Archive (Ismail Fawaz et al., 2019) Human - P
FordA UCR Archive (Ismail Fawaz et al., 2019) Sensor - P
InsectWingbeat UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ItalyPowerDemand UCR Archive (Ismail Fawaz et al., 2019) Power - P
ProximalPhalanxOutlineAgeGroup UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ACSF1 UCR Archive (Ismail Fawaz et al., 2019) Power - P
GunPoint UCR Archive (Ismail Fawaz et al., 2019) Human - P
RacketSports UCR Archive (Ismail Fawaz et al., 2019) Human - P
SmallKitchenAppliances UCR Archive (Ismail Fawaz et al., 2019) Power - P
ProximalPhalanxTW UCR Archive (Ismail Fawaz et al., 2019) Nature - P
DuckDuckGeese UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
PickupGestureWiimoteZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
EthanolLevel UCR Archive (Ismail Fawaz et al., 2019) Power - P
SpokenArabicDigits UCR Archive (Ismail Fawaz et al., 2019) Audio - P
SonyAIBORobotSurface1 UCR Archive (Ismail Fawaz et al., 2019) Facilities - P
HandOutlines UCR Archive (Ismail Fawaz et al., 2019) Human - P
PowerCons UCR Archive (Ismail Fawaz et al., 2019) Power - P
PhalangesOutlinesCorrect UCR Archive (Ismail Fawaz et al., 2019) Nature - P
BirdChicken UCR Archive (Ismail Fawaz et al., 2019) Nature - P
ToeSegmentation2 UCR Archive (Ismail Fawaz et al., 2019) Human - P
PigCVP UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P P P P
CricketY UCR Archive (Ismail Fawaz et al., 2019) Human - P
FingerMovements UCR Archive (Ismail Fawaz et al., 2019) Human - P
ElectricDevices UCR Archive (Ismail Fawaz et al., 2019) Power - P
InsectEPGRegularTrain UCR Archive (Ismail Fawaz et al., 2019) Nature - P
DodgerLoopGame UCR Archive (Ismail Fawaz et al., 2019) Traffic - P
Fungi UCR Archive (Ismail Fawaz et al., 2019) Nature - P
Symbols UCR Archive (Ismail Fawaz et al., 2019) Generated - P
MixedShapesRegularTrain UCR Archive (Ismail Fawaz et al., 2019) Generated - P
ArticularyWordRecognition UCR Archive (Ismail Fawaz et al., 2019) Human - P
UWaveGestureLibraryZ UCR Archive (Ismail Fawaz et al., 2019) Human - P
Epilepsy UCR Archive (Ismail Fawaz et al., 2019) Human - P
Healthcare200 UCR Archive (Ismail Fawaz et al., 2019) Healthcare - P
Weatherbench (Daily) Weather Bench Weather D P P possible P P P P P
Weatherbench (Hourly) Weather Bench Weather H P P possible P P P P P
Weatherbench (Weekly) Weather Bench Weather W P P possible P P P P P
Wiki hourly Wikipedia Metrics Web H P P
Wiki daily Wikipedia Metrics Web D P P P possible P P P P P
Wiki weekly Wikipedia Metrics Web W P P
Wiki monthly Wikipedia Metrics Web M P P
Global Temp Corrformer (Wu et al., 2023) Weather 10T T/T
TSMixup 10M Chronos (Ansari et al., 2024) Synthetic - P
KernelSynth 1M Chronos (Ansari et al., 2024) Synthetic - P
TaxiBJ Zhang et al., 2017 Traffic and weather 30T ZS
ERA5 Munoz-Sabater et al., 2021 Weather 3h P P
ERA5 Munoz-Sabater et al., 2021 Weather 12h P
ERA5 Munoz-Sabater et al., 2021 Weather D P
ERA5 Munoz-Sabater et al., 2021 Weather W P
ERA5 Munoz-Sabater et al., 2021 Weather M P
ERA5 Munoz-Sabater et al., 2021 Weather Q P
Finance (not published) Liu et al., 2025 (Sundial) Finance ? P
IoT (not published) Liu et al., 2025 (Sundial) IoT ? P
BOOM Cohen et al. (2025) (TOTO) Observational 10S ZS
BOOM Cohen et al. (2025) (TOTO) Observational T ZS
BOOM Cohen et al. (2025) (TOTO) Observational 5T ZS
BOOM Cohen et al. (2025) (TOTO) Observational 10T ZS
BOOM Cohen et al. (2025) (TOTO) Observational 30T ZS
BOOM Cohen et al. (2025) (TOTO) Observational H ZS
BOOM Cohen et al. (2025) (TOTO) Observational D